OpenStudy (anonymous):

Look at the formula for the area of a circle below. A = r2Which of the following equations can be used to solve for the radius (r)?

OpenStudy (solomonzelman):

the formula for the area of the circle is \(A=\pi {\rm r}^2\) (you left out \(\pi\) there)

OpenStudy (solomonzelman):

1) divide by \(\pi\) on both sides 2) square root both sides 3) normally when you do square roots you include \(\pm\) sign next to it, but in this case, since radius of the circle can be only positive (and not negative), you would not include the negative solution.